# Zhang, Liu, and Zhao

### v4

@@ 7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
where $I_{1}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$, $I_{2}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)$, and $I_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. With a alight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. Assumptions (B1) and (B2) hold in this case by construction. Assumption (B3) is justified since given a fixed surrogation index $I$, the surrogate function is independent of the reference point $\bar{\theta}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
8.3 Proofs of Theorems 3.6 and 3.7
In this section, we prove the convergence results in Theorems 3.6 and 3.7. As in Section 3.2.3, we suppose throughout this section that $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}$ of $F(\cdot,\xi)$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$.
@@
[…]
@@
[…]
@@
[…]
@@ 8.4 Nonasymptotic convergence result for general cases
Assumption (B5) has appeared in Proposition 8.10, and we state it here officially. Condition (P4) originating from Liu et al. (2022) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm. Note that Condition (P4) is verified for Case 1-Case 3 in Online Appendix 7. Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018). The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Online Appendix 8.5. It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
@@ 8.5 Verification of Assumption (B6)
For $\theta\in\mathrm{cl}(\mathcal{S}^{d})$, the claim holds clearly with $\theta^{d}=\theta$. For $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$, it suffices to consider the case with $n=1$, for which we denote $F(\theta)=\varphi(f(\theta))$, where $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ is piecewise linear and $\varphi$ is strongly convex. The polytope $\Theta$ can be partitioned into a finite number of $q$-polytopes, denoted by $\mathfrak{C}\coloneqq\{\mathcal{C}_{i}\}_{i=1}^{N}$, such that $f$ is linear on each of them and the intersection of any two polytopes in $\mathfrak{C}$ is their common face. Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$. Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\theta^{\prime}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\theta^{\prime}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.
@@ 9 Constrained SP with Covariate Information
    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators. This ensures feasibility of the ERM solution.
Following the excess risk decomposition in Section 2.2, it is straightforward to see that the PADR-based ERM model for the three special cases (C1) and (C2) is nonasymptotically consistent with the original SP problem (9).
@@ 9.1 Asymptotic consistency of constrained PADR-based ERM
    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (9.1).
By Assumptions (A2) and (A4), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$. We first provide the excess risk bound of the restricted ERM model (9) to the restricted SP problem (9.1).
@@
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (A4) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$. With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
@@
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
Proof 9.2
@@
where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$. The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $\Xi$,
$\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
With the excess risk bound in Proposition 9.1, and the probabilistic feasibility guarantee according to Luedtke and Ahmed (2008, Theorem 10), we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (9).

### v5

@@ 7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
where $I_{1}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$, $I_{2}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)$, and $I_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. With a slight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$. Assumptions (B1) and (B2) hold in this case by construction. Assumption (B3) is justified since given a fixed surrogation index $I$, the surrogate function is independent of the reference point $\bar{\theta}$.
@@
where $\left\lVert v\right\rVert=\left\lVert f(\theta,\xi)-f(\bar{\theta},\xi)\right\rVert\leq L_{f}\left\lVert\theta-\bar{\theta}\right\rVert$ by the Lipschitz continuity of $f$. By dividing the both sides by $\left\lVert\theta-\bar{\theta}\right\rVert^{2}$ and taking $\limsup$ as $\theta\rightarrow\bar{\theta}$ and $\theta\neq\bar{\theta}$, we obtain that $\widehat{\kappa}$ is finite.∎
@@
The assumption is naturally satisfied when the outer function $\varphi$ of $F$ is in Case 1 or Case 2. Specifically, from the proof of Claim 1 afterward, we see that in these two cases, $\Theta$ can be decomposed into finitely many convex regions, on each of which any d-stationary point $\theta^{d}\in\mathcal{S}^{d}$ is a minimizer of $F$ over that region. Since the number of such regions is finite, the function $F$ takes finitely many values on $\mathcal{S}^{d}$, verifying Assumption (EC1). Discussions on the finite number of d-stationary values for more general nonconvex problems with piecewise and composite structures can be seen in Cui and Pang (2018).
@@
The following lemma from Deng and Gao (2021) is rephrased in our context for the proof of Proposition 3.6.
Lemma 8.4 (Deng and Gao (2021), Theorem 3.2)
@@
where the third inequality follows from inequalities (3.8)-(3.13) in Mohri et al. (2018) that are part of the proof of Lemma 6.5, and the last inequality is derived by Lemma 6.3.
@@
8.3 Proofs of Theorems 3.7 and 3.8
In this section, we prove the convergence results in Theorems 3.7 and 3.8. As in Section 3.2.3, we suppose throughout this section that $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}$ of $F(\cdot,\xi)$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$.
@@
We next present the proof of Proposition 8.6. The overall scheme of the error-bound analysis follows the error-bound theory developed in Liu et al. (2022b) for the post-convergence analysis of the SMM algorithm. To justify the global error bound in the second statement of Proposition 8.6, we first give the error bounds of distance to $\mathcal{S}^{d}$ in terms of the residual $r_{d}$ in Proposition 8.8 and then relate $r_{d}$ with $r^{\varepsilon,\rho}$ in Proposition 8.10.
@@
A similar result to Proposition 8.10 is provided in Liu et al. (2022b, Lemma 4) under the assumption that the relevant surrogate function satisfies Condition (P4). With the surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot;\bar{\theta})$ associated with $r^{\varepsilon,\rho}(\bar{\theta})$ satisfying (P4), we can prove Proposition 8.10 following the lines of Liu et al. (2022b, Lemma 4), and thus the details are omitted here.
@@
Since $\lim_{t(\in\mathcal{K})\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$, we deduce $\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})=\theta_{\infty}$. Additionally, for a fixed index combination $\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})$, the surrogate function $\widehat{F}(\cdot;\cdot,\mathbf{I})$ is continuous on $\Theta\times\Theta$ by assumption, and thus by Shapiro et al. (2021b, Theorem 7.23), $\widehat{P}^{\rho}_{\mathbf{I}}(\cdot)$ is continuous at $\theta_{\infty}$. It follows from (8.11) that
@@ 8.4 Nonasymptotic convergence result for general cases
Assumption (B5) has appeared in Proposition 8.10, and we state it here officially. Condition (P4) originating from Liu et al. (2022a) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm. Note that Condition (P4) is verified for Case 1-Case 3 in Online Appendix 7. Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018a). The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Online Appendix 8.5. It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
@@ 8.5 Verification of Assumption (B6)
[…]
@@ 9 Constrained SP with Covariate Information
The optimization problem for learning the optimal decision rule can be formulated as
$\underset{\displaystyle\begin{subarray}{c}f\in\mathcal{F}\end{subarray}}{\mathrm{inf}}\quad\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (32)
@@
Constraints bring up the feasibility issue for data-driven decision rule methods. We show that in the following three special cases, probabilistic constraints can be equivalently converted into deterministic ones or embedded into the structure of PADR, with which we can construct the PADR-based ERM problem while retaining the feasibility of the obtained PADR:
@@
    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators. This ensures feasibility of the ERM solution.
Following the excess risk decomposition in Section 2.2, it is straightforward to see that the PADR-based ERM model for the three special cases (C1) and (C2) is nonasymptotically consistent with the original SP problem (32). For Case (C1), problems with finitely many scenarios can be addressed using the exact penalization formulation (37) developed in Section 9.2, and the feasibility of the solution is guaranteed by Lemma 9.7.
For the case with general constraints, it is difficult to convert them into deterministic ones, so we impose constraints only on finite scenarios in $\Xi$ and we formulate a restricted PADR-based ERM problem with an additional constant $\gamma>0$ as below,
$\underset{\displaystyle\begin{subarray}{c}f\in{\mathcal{H}}_{\mu}^{K}\end{subarray}}{\mathrm{min}}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})},{y^{s}})\hfil\hfil\hfil\hfil$ (33)
@@
Let ${\widetilde{\mathcal{H}}}^{K}_{\mu,\gamma}\coloneqq{\mathcal{H}}_{\mu}^{K}\bigcap\left\{f\in\mathcal{F}\mid\psi_{j}(f(x^{s}))+\gamma\leq 0,\,s\in[n],\,j\in[J]\right\}$ , and let $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ denote an optimal PADR to the constrained ERM problem (33). The idea of imposing the restricted constraints on finite scenarios follows the scenario approximation method in Calafiore and Campi (2005). This ensures that the obtained PADR $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ at a new scenario is feasible to the original constraint set with high probability according to Luedtke and Ahmed (2008, Theorem 10). The result is provided in the following proposition.
Proposition 9.1
Suppose that the support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$, and that $\psi_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L_{\psi}$-Lipschitz continuous for every $j\in[J]$. Then, with probability at least $1-\delta$ over the draw of $\Xi$, we have:
$\mathbb{P}_{X}\left[\max_{j\in[J]}\left\{\psi_{j}\left(f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}(X)\right)\right\}>0\right]<$ $\ \sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}$
$+\sqrt{\frac{\log(2n)}{n}}+\sqrt{\frac{2\log(1/\delta)}{n}}.$
Proof 9.2
Proof of Proposition 9.1 Denote $G(\theta,x)\coloneqq\max_{j\in[J]}\psi_{j}(f(\theta,x))$, where $\theta\in[-\mu,\mu]^{2dK(p+1)}$ is the parameter of PADR $f\in{\mathcal{H}}_{\mu}^{K}$. Since each $\psi_{j}$ for $j\in[J]$ is $L_{\psi}$-Lipschitz continuous, we have that for any $x\in\mathcal{X}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
$\left\lvert G(\theta_{1},x)-G(\theta_{2},x)\right\rvert$ $\leq L_{\psi}\left\lVert f(\theta_{1},x)-f(\theta_{2},x)\right\rVert$
$\leq L_{\psi}\sqrt{d}\left\lVert f(\theta_{1},x)-f(\theta_{2},x)\right\rVert_{\infty}$
$\leq 2L_{\psi}\sqrt{d(p\bar{X}^{2}+1)(p+1)}\cdot\left\lVert\theta_{1}-\theta_{2}\right\rVert_{\infty}.$
Define the sets
$\Theta^{\Xi}_{\gamma}\coloneqq\left\{\theta\in[-\mu,\mu]^{2dK(p+1)}\mid G(\theta,x^{s})\leq 0,\ s=1,\dots,n\right\},$
$\Theta_{\epsilon}\coloneqq\left\{\theta\in[-\mu,\mu]^{2dK(p+1)}\mid\mathbb{P}_{X}(G(\theta,X)\leq 0)\geq 1-\epsilon\right\}.$
By Luedtke and Ahmed (2008, Theorem 10), it follows that
$\mathbb{P}_{\Xi}\left(\Theta^{\Xi}_{\gamma}\subseteq\Theta_{\epsilon}\right)\geq 1-\left\lceil\frac{2}{\epsilon}\right\rceil\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)}e^{-n\epsilon^{2}/2}.$
With $\epsilon\in(0,1]$, let
$M\coloneqq\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)},\ \delta\coloneqq\left\lceil\frac{2}{\epsilon}\right\rceil Me^{-n\epsilon^{2}/2}\geq\frac{2}{\epsilon}Me^{-n\epsilon^{2}/2}.$
Without loss of generality, we can assume that $L_{\psi}\geq\gamma$, which leads to $M>2$ and $\log(\delta/(2M))<\log(1/4)<-1$. Therefore, we have that
$\frac{\delta}{2M}\geq\frac{1}{\epsilon}e^{-\frac{n}{2}\epsilon^{2}}>e^{(\log(\frac{\delta}{2M})+1)\frac{n}{2}\epsilon^{2}}\cdot e^{-\frac{n}{2}\epsilon^{2}}=e^{\log(\frac{\delta}{2M})\frac{n}{2}\epsilon^{2}},$
implying $\epsilon>\sqrt{2/n}$. Furthermore, denote $\delta^{\prime}\coloneqq Me^{-n\epsilon^{2}/2}$, it follows that with probability at least $1-\delta$ over $\Xi$, the solution $f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}$ satisfies:
$\mathbb{P}_{X}\left[\max_{j\in[J]}\left\{\psi_{j}\left(f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}(X)\right)\right\}>0\right]$
$\leq\sqrt{\frac{2}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)}/\delta^{\prime}\right)}$
$\leq\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta^{\prime})}{n}}$
$\leq\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta)}{n}}+\sqrt{\frac{2\log(2/\epsilon)}{n}}$
$<\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta)}{n}}+\sqrt{\frac{\log(2n)}{n}},$
which completes the proof. ∎
It can be observed that the out-of-sample feasibility of the PADR-based ERM solution exhibits the same order of complexity $\widetilde{\mathcal{O}}\left(\sqrt{\frac{4dK(p+1)}{n}}\right)$ in $n$ and $K$ as the generalization bound in the unconstrained case. Similar to the excess risk result for unconstrained problems, with a larger piece number $K$, model misspecification reduces but the likelihood of out-of-sample infeasibility simultaneously increases.
@@ 9.1 Asymptotic consistency of constrained PADR-based ERM
By strengthening the constraints by a constant parameter $\gamma_{0}>0$, we obtain a restricted SP problem as follows:
$\underset{\displaystyle\begin{subarray}{c}f\in\mathcal{F}\end{subarray}}{\mathrm{inf}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (34)
@@
for which we consider the following assumption, as an adaptation of Assumption (A4) for constrained problems:
@@
    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (34).
By Assumptions (A2) and (A4), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$. We first provide the excess risk bound of the restricted ERM model (33) to the restricted SP problem (34).
@@
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (A4) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$. With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
@@
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
Proof 9.4
Proof of Proposition 9.3. Denote by $f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ the optimal solution to the following problem
$\underset{\displaystyle\begin{subarray}{c}f\in{\mathcal{H}}_{\mu}^{K}\end{subarray}}{\mathrm{min}}\quad\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (35)
@@
Similar to analysis for unconstrained problems, we deduce the following decomposition:
@@
where the last inequality holds since $\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$. The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $\Xi$,
$\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
With the excess risk bound in Proposition 9.3, and the probabilistic feasibility guarantee according to Proposition 9.1, we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (32).